OpenStudy (anonymous):

Follow these instructions to make a geometric design using a compass and a straightedge. Then answer the question. SO CONFUSED :(

OpenStudy (anonymous):

1. Use a compass to draw a large circle. 2. Without changing the compass setting, place the point of the compass on the circle and make a mark that intersects the circle. 3. Move the point of the compass to the mark you made and make another mark that intersects the circle. 4. Continue in this way around the circle until you come back to the first mark. (Hint: The last mark might not coincide precisely with the first mark, but your drawing should look something like the sketch.) 5. Select one of the marks on the circle and use a straightedge to draw line segments connecting that mark to all the marks that are not next to it. 6. Repeat the process for each mark on the circle. How many line segments did you draw?

OpenStudy (anonymous):

I think it all comes out to 8 line segements?

OpenStudy (unklerhaukus):

something like this

jimthompson5910 (jim_thompson5910):

You're basically drawing diagonals of a hexagon there are n(n-3)/2 diagonals of any n-gon so there are n(n-3)/2 = 6(6-3)/2 = 6*3/2 = 9 diagonals total

OpenStudy (anonymous):

THANKS :)